OpenStudy (anonymous):

During his 2001 MVP season for the Seattle Mariners, Japanese baseball sensation Ichiro Suzuki hit 190 singles, 32 doubles, 10 triples, and 10 home runs in a total of 692 at-bats. The following table arranges these data in terms of the number of bases for a hit, counting 0 bases for the times when he did not get a hit, and 4 bases for home runs. Number of Bases 0 1 2 3 4 Number of Times 450 190 32 10 10 What was E(X), the expected number of bases for Ichiro at a typical at-bat in 2001? If necessary, use a fraction bar to enter a fraction. Reduce fractions to their lowest

OpenStudy (kropot72):

First set up a probability distribution table as below, where x is the number of bases. x 0 1 2 3 4 P(X = x) 450/692 190/692 32/692 10/692 10/692 \[E(X)=\sum_{}^{}xp(x)\] \[E(X)=0\times\frac{450}{692}+1\times\frac{190}{692}+2\times\frac{32}{692}+3\times\frac{10}{692}+4\times\frac{10}{692}=you\ can\ calculate\]

OpenStudy (anonymous):

Thanks!

OpenStudy (kropot72):

You're welcome :)